# Expander estimates for cubes

### v1

@@ 6. Minor arcs: first reductions
and hence that $a\geq 1$. Given the square-free parameter $w$ and $\alpha\in\mathfrak{m}$, any divisor $d$ of $w$ splits as $d=d_{q}d_{q}^{\prime}$ with $d_{q}^{\prime}=(d,q)$. Then $(d_{q},d_{q}^{\prime})=1$. Now let $\eta$ denote a positive number, and consider the sets
$\mathscr{E}$ =$=\{d|w:d\leq H,|G(\alpha d^{3},P/d,H/d)|\leq P^{1+\eta}d_{q}^{-1}\},$
$\mathscr{D}$ =$=\{d|w:d\leq H,|G(\alpha d^{3},P/d,H/d)|>P^{1+\eta}d_{q}^{-1}\}.$
@@
holds uniformly for all square-free $w$ composed of prime factors not exceeding $P$. It follows that
@@
We now apply Lemma 3.2 to bound $J$, and we use (2.3) together with Lemma 2.3 to control the Gauß sum. Then we conclude that
@@ 7. Endgame
#$\#\mathscr{D}^{\ddagger}(D,R,B)\ll(1+B)R^{2}D^{6}P^{\varepsilon-5/2}+DP^{\varepsilon-1/4}.$
@@
By (6.12), the set $\mathscr{D}^{\ddagger}(D,R,B)$ is empty unless $D\geq\frac{1}{2}P^{-1/4}$. Since $R^{3/2}D^{4}\leq P^{2-\eta}$, it follows that $\mathfrak{f}_{w}^{\ddagger}(D,R,B)\ll P^{\varepsilon-1/2}$. By (7.4), we then have $\mathfrak{f}_{w}^{\ddagger}\ll P^{\varepsilon-1/2}$. In view of Lemmas 6.3 and 7.1, we now have
@@
uniformly for $\alpha\in\mathfrak{m}$ and $w$ a square-free number free of prime factors exceeding $P$. Since the positive number $\eta$ is at our disposal, this argument establishes Theorem 1.1 for $\alpha\in\mathfrak{m}$.
This leaves the task to prove Lemma 7.2. To lighten notational burdens somewhat, we abbreviate #$\#\mathscr{D}^{\ddagger}(D,R,B)$ to $V$. With $D,R,B$ now fixed, let $V(A)$ denote the number of $d\in\mathscr{D}^{\ddagger}(D,R,B)$ where
@@ 8. Le coup de grâce
$h(3x^{2}+3xh+h^{2})=z_{1}-z_{2}$
while the constraints (8.4) imply
@@
Let $\mathscr{A}$ be a set of exponential density $\delta>0$, and write $\mathscr{A}(N)=\mathscr{A}\cap[1,N]$. Suppose that the real number $\eta$ satisfies $0<\eta<\delta$. Then, there are infinitely many natural numbers $N$ where $A_{N}=\#\mathscr{A}(N)\geq N^{\eta}$. With such an $N$, we apply Theorem 8.2 with $\mathscr{Z}=\mathscr{A}(N)$. Then $\Theta(N)$ counts certain elements of $\mathscr{B}_{3}$ (defined in (1.1)), and all these elements do not exceed $(2P)^{3}+N\leq 9P^{3}$, at least when $N$ is large. It follows that
#$\#\mathscr{B}_{3}(9P^{3})\gg P^{-\varepsilon}\min(P^{3},PA_{N})\gg P^{-\varepsilon}\min(P^{3},P^{1+5\eta/2}).$
It follows that $\delta_{3}\geq\min(1,\frac{1}{3}+\frac{5}{6}\eta)$, for all $\eta<\delta$. This establishes Theorem 1.1. ∎

### v2

@@ 6. Minor arcs: first reductions
and hence that $a\geq 1$. Given the parameter $w$ and $\alpha\in\mathfrak{m}$, any square-free divisor $d$ of $w$ splits as $d=d_{q}d_{q}^{\prime}$ with $d_{q}^{\prime}=(d,q)$. Then $(d_{q},d_{q}^{\prime})=1$. Now let $\eta$ denote a positive number, and consider the sets
$\mathscr{E}$ $=\{d|w:\mu(d)^{2}=1,d\leq H,|G(\alpha d^{3},P/d,H/d)|\leq P^{1+\eta}d_{q}^{-1}\},$
$\mathscr{D}$ $=\{d|w:\mu(d)^{2}=1,d\leq H,|G(\alpha d^{3},P/d,H/d)|>P^{1+\eta}d_{q}^{-1}\}.$
@@
holds uniformly for all $w$ composed of prime factors not exceeding $P$. It follows that
@@
We now apply Lemma 3.2 to bound $J$, and Lemma 2.4 to control the Gauß sum. Then we conclude that
@@ 7. Endgame
$\mathscr{D}^{\ddagger}(D,R,B)\ll(1+B)R^{2}D^{6}P^{\varepsilon-5/2}+DP^{\varepsilon-1/4}.$
@@
By (6.12), the set $\mathscr{D}^{\ddagger}(D,R,B)$ is empty unless $D\geq\frac{1}{2}P^{1/4}$. Since $R^{3/2}D^{4}\leq P^{2-\eta}$, it follows that $\mathfrak{f}_{w}^{\ddagger}(D,R,B)\ll P^{\varepsilon-1/2}$. By (7.4), we then have $\mathfrak{f}_{w}^{\ddagger}\ll P^{\varepsilon-1/2}$. In view of Lemmas 6.3 and 7.1, we now have
@@
uniformly for $\alpha\in\mathfrak{m}$ and $w$ a number free of prime factors exceeding $P$. Since the positive number $\eta$ is at our disposal, this argument establishes Theorem 1.2 for $\alpha\in\mathfrak{m}$.
This leaves the task to prove Lemma 7.2. To lighten notational burdens somewhat, we abbreviate $\mathscr{D}^{\ddagger}(D,R,B)$ to $V$. With $D,R,B$ now fixed, let $V(A)$ denote the number of $d\in\mathscr{D}^{\ddagger}(D,R,B)$ where
@@ 8. Le coup de grâce
$h(3x^{2}+3xh+h^{2})=z_{2}-z_{1}.$
By (8.4), we first see that $h\geq 1$, and then that $3x^{2}+3xh+h^{2}>P^{2}$. Recalling now that $N=P^{5/2}$, we see that the constraints (8.4) together with the preceding equation imply conditions
@@
Let $\mathscr{A}$ be a set of exponential density $\delta>0$, and write $\mathscr{A}(N)=\mathscr{A}\cap[1,N]$. Suppose that the real number $\delta^{\prime}$ satisfies $0<\delta^{\prime}<\delta$. Then, there are infinitely many natural numbers $N$ where $A_{N}=\#\mathscr{A}(N)\geq N^{\delta^{\prime}}$. With such an $N$, we apply Theorem 8.2 with $\mathscr{Z}=\mathscr{A}(N)$. Then $\Theta(N)$ counts certain elements of $\mathscr{B}_{3}$ (defined in (1.1)), and all these elements do not exceed $(2P)^{3}+N\leq 9P^{3}$, at least when $N$ is large. It follows that
$\mathscr{B}_{3}(9P^{3})\gg P^{-\varepsilon}\min(P^{3},PA_{N})\gg P^{-\varepsilon}\min(P^{3},P^{1+5\delta^{\prime}/2}).$
It follows that $\delta_{3}\geq\min(1,\frac{1}{3}+\frac{5}{6}\delta^{\prime})$, for all $\delta^{\prime}<\delta$. This establishes Theorem 1.1. ∎